1. Using the definition C = Q/V, show that for a SERIES connection of capacitors:



1/C1/C1+1/C2. What does this equation mean? What happens to C as more and more capacitors are connected in series? 2. Using the definition C = QV, show that for a PARALLEL connection of capacitors: C = C1 + C2. What does this equation mean? What happens to C as more and more capacitors are connected in parallel?

1
Expert's answer
2022-03-28T11:02:23-0400

1. For series connection

"Q_1=Q_2=Q,\quad V_1+V_2=V"

Hence

"V=V_1+V_2=\frac{Q_1}{C_1}+\frac{Q_1}{C_1}=\frac{Q}{C}"

Finally

"\frac{1}{C_1}+\frac{1}{C_1}=\frac{1}{C}"

2. For parallel connection

"Q_1+Q_2=Q,\quad V_1=V_2=V"

Hence

"C_1V_1+C_2V_2=CV"

Finally

"C_1+C_2=C"